­­­­­­LHC 2011 - latest news

Week 40: Machine coordinators: Ralph Assmann and Barbara Holzer

Goal of the week: Luminosity production




Link to report from morning meeting



8:30 meeting

 Sunday 9th October

  • 4:58 programmed dump after 17:17 hours in stable beams. : integrated luminosity Atlas/CMS/LHCb: 115/120/21.3 pb-1; Alice: 58.5 nb-1
  • 6:20 injecting physics beams
  • 6:27 Extraction kicker problem for B1 in the SPS.
  • 6:54 beams dumped (B2 already fully injected) as the SPS cannot deliver beams before 9:00
  • 9:11 beam back
  • 11:11 stable beams Fill #2197 Initial luminosity: 3.42E33 cm-2 s-1! Slightly higher bunch intensities (1.35E11p) and again an excellent beam from the injectors!
  • 18:02 Beams dumped. Trip of S56, PIC communication lost.

9:00 meeting

 Saturday 8th October

  • 01:44 Injection for physics. Kicker problem solved. Injection efficiency is affected by a BPM interlock in SPS-LSS6 for B1. SPS problems. Injected beam dumped.
  • 03:00 Injection for physics again.
  • 03:51 Ramp.
  • 04:33 Stable beams. Fill #2194. Initial luminosity: 3.2e33.
  • 05:10 Beams dumped by trip of sector 56. MP3 investigating. Integrated lumi: 7pb-1. Ramp down and recycle.
  • 07:41 Access to point 6 (QPS problem) and point 2 (ALICE).
  • 08:50 Precycling.
  • 10:00 injecting probe beams
  • 11:41 Stable Beams Fill #2195. Initial luminosity: 3.3e33.

9:00 meeting

  Friday 7th October

  • 02:30 preparation of 25ns physics, 60 bunches/beam
  • 05:23 stable beams Fill #2186 (emittance in collision ~3.3um, bunch length 01.32/1.38ns, lifetime 125 hours) peak lumi ~0.06e33 cm-2 s-1
  • 08:16 dumped by RF; integrated lumi 0.6 nb-1.
  • 09:14 BIC problem in TI8 and CMS recovering their tracker
  • 10:09 Abort gap cleaning.
  • 16:29 End of abort gap cleaning. Dump. Rampdown and precycle.
  • 16:58 RF lines 7B1, 7B2 and 8B2 have an error on the LL.
  • 17:59 RF back. Checks for the new XPOC release.
  • 18:32 Injection. Beam 2 injection problems. Transfer line steering. MKI8 faults. TI2 BPM LSS6 interlocks.
  • 22:20 Machine filled. Start ramp.
  • 22:29 Beams dumped. Trips 2 RF modules, arcs detected. Ramp down and precycle.
  • 23:28 Injection. Injection problems. SPS injection oscillations. MKI8 faults.
  • 23:55 Called kicker expert. The problem with the MKI is that now there is a shift between the prepulse and the abort gap keeper.  Expert comes to CCC to correct the shift. RF checked that all OK from their side.

8:30 meeting
report on 'general power off'

 Thursday 6th October

  • 07:43 Accidental switch off LHC magnets including experiments
  • Accesses in all IPs except IP3 (EL, CV, RF, ADT, vacuum, optical fibers, EPC, diamond BLM, vacuum, QPS and CNGS)
  • 12:02 Patrol in Pt2 (end of vacuum intervention on TED)
  • LHCb polarity changed
  • 14:30 beam back for 25ns MD
  • till 02:30 very successful 25ns MD (injection train of 288 bunches B2, 144 bunches B1; 732 bunches circulating B2 with 5h lifetime; ecloud, vacuum and cryo activity)

8:30 meeting

 Wed 5th October

  • 05:31 Beam dump due to QPS trigger in sector 78. Quench heaters discharged in MB10.L8.
  • 07:28 Quench heaters charged ok again in cell B10L8. Was a EN/EL circuit breaker (cell 32, EOJ106/78.32R7). nr. Q3.
  • 08:29 Cryo back.
  • 08:20 RF tests without beam. Check for problem of debunching observed during fill #2182.
  • 09:51 Vacuum leak in TI2 TED bellows. Access breaks patrol. Initial estimate for repair and recovery: 48h.
  • 15:16 Open LHC for access. ELQA access in P8. Planning for beam 2 program over night: abort gap cleaning and 25ns setup.
  • 18:09 Vacuum group identifies correct spare part. Recovery time revised to Thursday morning. TED repair starting. Beam 2 program for the night cancelled. Plan TI8 stability checks overnight (during pump down of TI2).
  • 18:32 Starting the PNO.a7 on RQ5.L8 and RQ6.L8.
  • 19:03 Going through access list to check who wants to go into LHC over night and during early morning.
  • 20:10 Access in P8 (vac expert + RP) for preventive check of the TED bellows in TI8.

8:30 meeting

 Tuesday 4 October

  • 12:12 Fill #2181 dumped.
  • RF testing new tuner firmware in preparation for the 25ns MD on Thursday
  • 14:00 RF problem (rollback to previous firmware version)
  • 14:50 BLM problem
  • abort gap cleaning MD moved to Friday afternoon
  • 17:53 injection for physics
  • 19:48 start of ramp
  • 20:36 stable beams Fill #2182 (1.8E14 protons, peak lumi ~3.2e33 cm-2 s-1)

8:30 meeting

Mon 3rd October

  • 09:25 Operator dump fill #2178, integrated lumi 107/108/18 nb-1. Trip of RQ4.R6 (fire heaters discharged) just after the beam dump but the QPS gave no PM data.
  • 10:27 Try to reset Q4.R6. Not possible. Access needed.
  • 11:03 Prepare access to UA area in point 6.
  • 12:17 Prepare tunnel access in point 6.
  • 14:27 Lost cryo conditions in the MSR6.
  • 14:53 Q4.R6 problem solved (~5h impact). The firing of the heaters has somehow affected the current lead instrumentation. Could not reset the QPS of the Q4. Was issue with the ELQA proximity box, on the current lead side. The fault was cleared by de-plugging and re-connecting the cable but it is not clear what the source was. The QPS and ELQA teams are still working to try and understand more in details what happened.
  • 15:36 Pre-cycle. Some RF cavity studies (Philippe). Some BLM thresholds changed.
  • 16:55 Injection.
  • 17:28 Optimization of transfer line is required (~1h impact).
  • 18:40 Injector beam optimization (~0.7h impact).
  • 19:36 Filling complete.
  • 19:45 Ramp.
  • 20:34 Stable beams. Fill #2180. Peak lumi ~3.2e33 cm-2 s-1.
  • 20:48 Beams dumped due to high losses in IR2 after vacuum spike.
  • 21:41 Injection.
  • 23:57 Stable beams. Fill #2181. Peak lumi ~3.2e33 cm-2 s-1.

8:30 meeting